OpenStudy (anonymous):

Find all solutions to the equation. cos2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

i dont understand

OpenStudy (zehanz):

Do you know a formula for cos2x?

OpenStudy (zehanz):

Like cos2x = 2cos²x-1 ?

OpenStudy (anonymous):

no but i know cosx= (1/secx)

OpenStudy (zehanz):

If you replace cos2x with 2cos²x-1, you have a quadratic equation, with cosx as variable.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

error, my bad. cos^(2)x + 2 cos x + 1 = 0

OpenStudy (zehanz):

OK, so you already have a quadratic equation in cosx: \[(\cos x)^2 + 2\cos x +1 = 0\] If you like you could temporarily set p=cosx to get:\[p^2+2p+1=0\]Could you try to solve that one?

OpenStudy (anonymous):

im stuck at p^(3)=-1/2

OpenStudy (zehanz):

How do you get p^3 ?

OpenStudy (anonymous):

@Lauren01 is it cos^2 x or cos 2x

OpenStudy (zehanz):

This is the form you can use: (p+...)(p+...)=0. You must have done that before, I guess. Just look for the numbers on the dots: their sum is 2, their product is 1. They are not difficult to find...

OpenStudy (zehanz):

@Yahoo! It is cos²x, @Lauren01 already mentioned...

OpenStudy (anonymous):

oh...Sorry ...Did nt See !

OpenStudy (zehanz):

We're waiting for Lauren01 to make a move.. ;)

OpenStudy (anonymous):

sorry i got distracted by the office. its pie. because x=-1 is on Pie. (unit circle)

OpenStudy (anonymous):

(p+1)(p+1)=0 cosx+1=0 cosx=-1 using the unit cirle x=-1 is on Pie.

OpenStudy (zehanz):

Well done! BTW, aren't you supposed to give more answers, like x=pi+2k*pi, or do you have to solve only for x in [0,2pi]?

OpenStudy (anonymous):

only one is fine

OpenStudy (zehanz):

OK, next prob!

OpenStudy (anonymous):

lol do you wanna do all my homework? lmao

OpenStudy (zehanz):

Now you do your homework, then I'll only help you along...

OpenStudy (anonymous):

lol goood choice